# Uncertainty Quantification for Prior-Data Fitted Networks using Martingale Posteriors

### v2

@@ B.1 Estimation of $\rho$
We implemented the estimation of $\rho$ for the Gaussian copula update as follows: If the computational budget is limited, we follow the routine suggested by Fong et al., (2023) and maximize the prequential log score $\sum_{i=1}^{n}\log p_{i-1}(y_{i}|\bm{x}_{i})$ for $\rho$ (an efficient method is implemented in the MP module by the same authors). If the dataset size is limited or if one likes to separate tuning from training, one could potentially draw a new and larger set of density values from TabPFN and independently estimate using this “tuning set”. In practice, we found the training data to be sufficient for good estimation of $\rho$. We also reuse this estimator for all predictive steps in our Martingale Posterior computation and have not found any significant improvement by re-estimating the correlation or manually tuning via cross-validation.
Figure 7: Comparison of the posterior from our method (blue) and the one obtained by predictive resampling (red) for different shape parameters $\gamma$ (columns) of the Gamma distribution.
B.2 Additional Experiments
B.2.1 Unconditional Quantile Posterior
We here provide another experiment where we analyze the ability of our approach to estimate the posterior of an extreme quantile of a skewed distribution. For this, we simulate a Gamma distribution with different shape parameters $\gamma\in\{1,2,4\}$, inducing varying left-skewness. We then task PFN to estimate the 99%-quantile (the function $T$) and use our approach to compute the posterior uncertainty for PFN’s estimate. In this experiment, we use $B=1,000$ replications and $N=10,000$. To make the task of quantile estimation more challenging, we use a relatively small dataset size of $n=25$. To evaluate the performance, we compare the distribution against the true value and a posterior estimate by the predictive resampling approach from Fong et al., (2023) that does not have access to the PFN.
Results
An exemplary result for the posteriors for different shape values is visualized in Figure 7, showing the general trend of the results. The predictive resampling methods, which do not have access to the simulated data from TabPFN, usually result in a much more concentrated and bimodal posterior. This, however, comes at the cost of not always covering the true value. In contrast, the posterior for our method is much wider, thereby always covering the true value independent of the shape parameter.
B.2.2 Conditional Quantile Coverages
@@ Results
The data is visualized in Figure 10 in Appendix B. The resulting coverages are shown in Table 1. Our method provides (close to) nominal coverage for many combinations of $x$ and $\alpha$ but is less accurate for $x=0.2$ and extremer quantiles in general. This can be explained by the fact that the data at $x=0.2$ has almost no variation, while little data is available for extreme values, making these cases very challenging given only 100 data points.
Boostrap Comparison
@@
Table 2: Results for different values of $x$ and $\alpha$.
B.2.3 Convergence Speed: Additional Results
Figures 8 and 9 depict the convergence speeds for the analysis in Section 4.2 for the two other learning rate schedules, type 1 and type 2, respectively. In this example, all results look qualitatively very similar.
@@ B.2.3 Convergence Speed: Additional Results
B.2.4 Benchmarks: Additional Results
Tables 3 and 4 summarize the results of our benchmark for $\alpha=0.5$ and $\alpha=0.9$, respectively for $n=25$. While both methods often have very close to nominal coverage or provide larger than nominal coverage, the bootstrap method fails in one case to provide a CI with close to nominal coverage. Despite having larger than nominal coverage, our MP approach consistently yields shorter or equally large CIs while often being twice as fast as the bootstrap.
@@ B.2.4 Benchmarks: Additional Results
B.2.5 Benchmarks: Comparison Using Fixed Time Budget
@@
Table 5: Coverage and interval lengths for MP and Bootstrap methods.
B.2.6 Data Generation for Diffusion Process
For the diffusion process data in Section 4.3, we use the following data-generating process:
@@ B.2.6 Data Generation for Diffusion Process
B.2.7 Data Generation for Quantile Regression
@@
B.2.8 Prompt and Details for LLM Experiment
The prompt used for the experiment in our discussion is given as follows:
@@ B.2.8 Prompt and Details for LLM Experiment
B.3 Computational Environment
All computations were performed on a user PC with Intel(R) Core(TM) i7-8665U CPU @ 1.90GHz, 8 cores, 16 GB RAM, using Python 3.8, R 4.2.1, and TensorFlow 2.10.0. Run times of each experiment do not exceed 24 hours.
All experiments that use TabPFN are done with tabpfn==2.0.5. GPT 4o and GPT o3-mini-high are called via OpenAI’s UI. DeepSeek R1 is called from its UI at www.deepseek.com. All other methods are implemented from scratch by the authors.
Appendix C Further Discussion
Another open problem is joint posterior inference for a collection of parameters $\theta=\{\theta(x)\colon x\in\mathcal{X}\}$. Although our proposed inference procedure can be repeated for many values of $x$, the resulting posteriors are disconnected. For obtaining a full joint posterior $\Pi(\theta|\mathcal{D}_{n})$, the distribution of the features $x_{1:n}$ can no longer be ignored. Fong et al., (2023) proposed a general joint update of the PPDs for all values of $x$ simultaneously. However, this general update is intractable, and the heuristic simplifications proposed by Fong et al., (2023) are neither particularly simple nor theoretically justified. There are several potential ways forward. A simple practical solution would be to specify a joint distribution that combines the individual posteriors $P_{\infty,x_{1}},\dots,P_{\infty,x_{K}}$ in a plausible way; for example, using a multivariate Gaussian copula with covariance kernel depending on the distance between values $x_{i}\neq x_{j}$. We expect such a heuristic correction to work reasonably well in many applications. A more sophisticated alternative was recently proposed by Huk et al., (2024) and involves nonparametric estimation of the implicit dependence between PPDs by a nonparametric vine copula.

### v3

@@ B.1 Estimation of $\rho$
We implemented the estimation of $\rho$ for the Gaussian copula update as follows: If the computational budget is limited, we follow the routine suggested by Fong et al. (2023) and maximize the prequential log score $\sum_{i=1}^{n}\log p_{i-1}(y_{i}|\bm{x}_{i})$ for $\rho$ (an efficient method is implemented in the MP module by the same authors). If the dataset size is limited or if one likes to separate tuning from training, one could potentially draw a new and larger set of density values from TabPFN and independently estimate using this “tuning set”. In practice, we found the training data to be sufficient for good estimation of $\rho$. We also reuse this estimator for all predictive steps in our Martingale Posterior computation and have not found any significant improvement by re-estimating the correlation or manually tuning via cross-validation.
B.2 Additional Simulation Results
The details for the conditional coverages for all simulation settings of the study described in Section 4.3 can be found in Tables 3, 4 and 5
Table 3: Detailed simulation results for $n=30$ (coverage with interval length in brackets).
$\varphi$ $x$ Optimal Fong et al. Ours (Default) Ours (Type-1) Ours (Type-2) Pure TabPFN
1 -2.50 0.716 (2.113) 0.963 (2.588) 0.934 (2.219) 0.942 (2.254) 0.942 (2.254) 0.998 (3.152)
-1.94 0.947 (1.990) 0.947 (2.588) 0.943 (2.395) 0.949 (2.430) 0.949 (2.430) 0.945 (2.007)
-1.39 0.961 (2.148) 0.990 (2.588) 0.974 (2.307) 0.970 (2.307) 0.970 (2.307) 1.000 (5.846)
-0.83 0.960 (2.113) 0.956 (2.588) 0.931 (2.254) 0.945 (2.307) 0.945 (2.307) 0.985 (2.641)
-0.28 0.920 (1.831) 0.986 (2.571) 0.978 (2.377) 0.977 (2.395) 0.977 (2.395) 0.910 (1.902)
0.28 0.845 (1.937) 0.965 (2.571) 0.959 (2.342) 0.963 (2.360) 0.963 (2.360) 0.920 (2.148)
0.83 0.938 (1.831) 0.943 (2.571) 0.861 (2.360) 0.855 (2.360) 0.855 (2.360) 0.902 (2.024)
1.39 0.925 (1.867) 0.989 (2.571) 0.976 (2.272) 0.976 (2.289) 0.976 (2.289) 0.788 (2.258)
1.94 0.941 (1.955) 0.604 (2.553) 0.589 (2.324) 0.604 (2.324) 0.604 (2.324) 0.564 (2.208)
2.50 0.910 (2.095) 0.939 (2.553) 0.921 (2.272) 0.938 (2.324) 0.938 (2.324) 0.967 (2.623)
3 -2.50 0.285 (2.787) 0.906 (3.612) 0.906 (3.592) 0.906 (3.592) 0.906 (3.592) 0.998 (7.830)
-1.94 0.287 (2.787) 0.910 (3.612) 0.896 (3.651) 0.896 (3.690) 0.896 (3.690) 0.947 (3.442)
-1.39 0.987 (2.945) 0.999 (3.612) 0.998 (3.494) 0.999 (3.553) 0.999 (3.553) 1.000 (4.072)
-0.83 0.869 (2.905) 0.989 (3.455) 0.995 (3.533) 0.995 (3.553) 0.995 (3.553) 0.936 (2.280)
-0.28 0.893 (2.552) 0.441 (2.925) 0.900 (3.514) 0.900 (3.533) 0.900 (3.533) 0.697 (3.600)
0.28 0.774 (2.768) 0.001 (1.963) 0.748 (3.533) 0.748 (3.553) 0.748 (3.553) 0.573 (3.090)
0.83 0.731 (2.630) 0.087 (1.924) 0.989 (3.514) 0.987 (3.592) 0.987 (3.592) 0.973 (7.290)
1.39 0.993 (3.239) 0.782 (2.768) 0.995 (3.416) 0.995 (3.416) 0.995 (3.416) 0.960 (2.878)
1.94 0.893 (3.062) 0.856 (3.416) 0.942 (3.514) 0.942 (3.553) 0.942 (3.553) 0.794 (2.696)
2.50 0.651 (2.729) 0.500 (3.612) 0.439 (3.455) 0.439 (3.475) 0.439 (3.475) 0.771 (4.372)
10 -2.50 0.977 (5.679) 1.000 (6.659) 1.000 (6.017) 1.000 (6.085) 1.000 (6.085) 0.970 (6.390)
-1.94 0.970 (5.679) 1.000 (6.659) 0.998 (5.949) 0.998 (5.983) 0.998 (5.983) 0.948 (5.401)
-1.39 1.000 (5.713) 1.000 (6.659) 1.000 (6.051) 1.000 (6.118) 1.000 (6.118) 1.000 (6.645)
-0.83 1.000 (5.442) 1.000 (6.659) 1.000 (5.882) 1.000 (5.882) 1.000 (5.882) 0.982 (5.352)
-0.28 0.865 (5.307) 0.996 (6.659) 0.936 (6.118) 0.967 (6.220) 0.967 (6.220) 0.773 (5.792)
0.28 0.974 (5.476) 1.000 (6.659) 1.000 (6.186) 1.000 (6.186) 1.000 (6.186) 0.995 (6.247)
0.83 1.000 (5.713) 1.000 (6.625) 1.000 (6.118) 1.000 (6.220) 1.000 (6.220) 0.969 (6.250)
1.39 0.678 (5.442) 0.998 (6.625) 0.968 (5.611) 0.968 (5.611) 0.968 (5.611) 0.683 (5.014)
1.94 0.431 (5.645) 0.953 (6.558) 0.842 (5.848) 0.842 (5.780) 0.842 (5.780) 0.426 (5.385)
2.50 0.340 (5.611) 0.922 (6.592) 0.831 (6.186) 0.865 (6.220) 0.865 (6.220) 0.668 (5.842)
Table 4: Detailed simulation results for $n=100$ (coverage with interval length in brackets).
$\varphi$ $x$ Optimal Fong et al. Ours (Default) Ours (Type-1) Ours (Type-2) Pure TabPFN
1 -2.50 0.959 (2.010) 0.998 (3.134) 0.945 (1.912) 0.945 (1.912) 0.945 (1.912) 0.964 (2.247)
-1.94 0.902 (1.853) 0.786 (3.075) 0.801 (1.675) 0.801 (1.675) 0.801 (1.675) 0.766 (1.261)
-1.39 0.922 (1.754) 0.990 (2.937) 0.890 (1.596) 0.890 (1.596) 0.890 (1.596) 0.927 (2.148)
-0.83 0.898 (1.794) 0.977 (2.621) 0.863 (1.557) 0.863 (1.557) 0.863 (1.557) 0.917 (1.872)
-0.28 0.888 (1.715) 0.475 (2.247) 0.892 (1.675) 0.892 (1.675) 0.892 (1.675) 0.770 (1.202)
0.28 0.913 (1.715) 0.923 (2.050) 0.902 (1.794) 0.902 (1.794) 0.902 (1.794) 0.922 (1.899)
0.83 0.924 (1.754) 0.379 (2.129) 0.935 (1.813) 0.935 (1.813) 0.935 (1.813) 0.941 (1.907)
1.39 0.951 (1.931) 0.983 (2.385) 0.986 (2.385) 0.986 (2.385) 0.986 (2.385) 0.924 (2.060)
1.94 0.900 (1.774) 0.986 (2.818) 0.856 (1.636) 0.856 (1.636) 0.856 (1.636) 0.907 (1.895)
2.50 0.947 (2.148) 0.997 (3.075) 0.932 (1.872) 0.932 (1.872) 0.932 (1.872) 0.991 (2.591)
3 -2.50 0.976 (2.243) 0.998 (3.602) 0.745 (2.786) 0.745 (2.786) 0.745 (2.786) 0.993 (5.536)
-1.94 0.902 (2.039) 0.950 (3.579) 0.978 (2.696) 0.978 (2.696) 0.978 (2.696) 0.998 (6.488)
-1.39 0.945 (1.948) 0.861 (3.579) 0.935 (2.673) 0.935 (2.673) 0.935 (2.673) 0.923 (2.512)
-0.83 0.905 (1.948) 0.998 (3.602) 0.985 (2.900) 0.985 (2.900) 0.985 (2.900) 0.976 (3.209)
-0.28 0.921 (1.880) 0.885 (3.443) 0.992 (3.013) 0.992 (3.013) 0.992 (3.013) 0.860 (2.647)
0.28 0.937 (1.835) 0.980 (3.239) 0.995 (2.786) 0.995 (2.786) 0.995 (2.786) 0.876 (2.466)
0.83 0.902 (1.903) 0.970 (3.579) 0.996 (2.990) 0.996 (2.990) 0.996 (2.990) 0.748 (2.433)
1.39 0.913 (1.926) 0.952 (3.602) 0.981 (2.786) 0.981 (2.786) 0.981 (2.786) 0.938 (3.049)
1.94 0.926 (2.152) 0.998 (3.602) 0.850 (2.718) 0.850 (2.718) 0.850 (2.718) 0.878 (2.461)
2.50 0.979 (2.288) 0.997 (3.602) 0.997 (2.877) 0.997 (2.877) 0.997 (2.877) 0.929 (2.588)
10 -2.50 0.980 (3.219) 0.004 (1.048) 0.999 (6.214) 0.999 (6.214) 0.999 (6.214) 0.996 (4.901)
-1.94 0.998 (3.294) 0.318 (1.048) 1.000 (6.289) 1.000 (6.289) 1.000 (6.289) 1.000 (7.105)
-1.39 0.992 (3.294) 0.006 (1.048) 1.000 (6.476) 1.000 (6.476) 1.000 (6.476) 0.999 (6.446)
-0.83 0.998 (3.107) 0.009 (1.048) 1.000 (6.214) 1.000 (6.214) 1.000 (6.214) 0.997 (6.578)
-0.28 0.997 (3.069) 0.004 (1.048) 1.000 (6.363) 1.000 (6.363) 1.000 (6.363) 0.992 (5.660)
0.28 0.996 (3.032) 0.001 (1.048) 0.999 (6.588) 0.999 (6.625) 0.999 (6.625) 0.950 (5.060)
0.83 0.991 (3.107) 0.005 (1.048) 1.000 (6.326) 1.000 (6.326) 1.000 (6.326) 0.943 (5.035)
1.39 0.997 (3.069) 0.094 (1.048) 1.000 (6.251) 1.000 (6.251) 1.000 (6.251) 0.969 (5.869)
1.94 0.999 (3.219) 0.038 (1.048) 1.000 (5.802) 1.000 (5.802) 1.000 (5.802) 0.802 (5.548)
2.50 0.924 (3.294) 0.000 (1.048) 0.995 (6.513) 0.995 (6.513) 0.995 (6.513) 0.701 (5.039)
Table 5: Detailed simulation results for $n=300$ (coverage with interval length in brackets).
$\varphi$ $x$ Optimal Fong et al. Ours (Default) Ours (Type-1) Ours (Type-2) Pure TabPFN
1 -2.50 0.954 (2.020) 0.935 (1.827) 0.926 (1.770) 0.926 (1.770) 0.926 (1.770) 0.982 (2.501)
-1.94 0.922 (1.866) 0.892 (1.770) 0.884 (1.770) 0.884 (1.770) 0.884 (1.770) 0.869 (1.520)
-1.39 0.894 (1.674) 0.613 (1.674) 0.879 (1.577) 0.879 (1.577) 0.879 (1.577) 0.861 (1.789)
-0.83 0.898 (1.674) 0.443 (1.577) 0.915 (1.750) 0.915 (1.750) 0.915 (1.750) 0.907 (1.712)
-0.28 0.907 (1.674) 0.756 (1.539) 0.835 (1.404) 0.835 (1.404) 0.835 (1.404) 0.698 (1.366)
0.28 0.903 (1.674) 0.502 (1.481) 0.897 (1.635) 0.897 (1.635) 0.897 (1.635) 0.932 (1.904)
0.83 0.899 (1.674) 0.793 (1.423) 0.951 (2.231) 0.951 (2.231) 0.951 (2.231) 0.975 (2.526)
1.39 0.898 (1.693) 0.780 (1.385) 0.954 (2.154) 0.954 (2.154) 0.954 (2.154) 0.961 (2.087)
1.94 0.922 (1.750) 0.101 (1.347) 0.932 (1.847) 0.932 (1.847) 0.932 (1.847) 0.957 (1.928)
2.50 0.920 (1.943) 0.001 (1.327) 0.909 (2.001) 0.909 (2.001) 0.909 (2.001) 0.783 (1.676)
3 -2.50 0.924 (2.113) 0.860 (3.090) 0.960 (2.681) 0.960 (2.681) 0.960 (2.681) 0.977 (3.849)
-1.94 0.927 (1.818) 0.945 (3.090) 0.776 (2.045) 0.776 (2.045) 0.776 (2.045) 0.950 (3.560)
-1.39 0.916 (1.727) 0.864 (3.090) 0.905 (1.863) 0.905 (1.863) 0.905 (1.863) 0.886 (2.920)
-0.83 0.915 (1.727) 0.950 (3.090) 0.948 (2.045) 0.948 (2.045) 0.948 (2.045) 0.934 (2.777)
-0.28 0.911 (1.727) 0.967 (3.090) 0.967 (2.181) 0.967 (2.181) 0.967 (2.181) 0.890 (2.365)
0.28 0.904 (1.704) 0.955 (3.090) 0.962 (2.158) 0.962 (2.158) 0.962 (2.158) 0.875 (2.817)
0.83 0.887 (1.749) 0.998 (3.090) 0.922 (2.158) 0.922 (2.158) 0.922 (2.158) 0.837 (2.542)
1.39 0.907 (1.727) 0.996 (3.090) 0.985 (2.499) 0.985 (2.499) 0.985 (2.499) 0.926 (2.900)
1.94 0.925 (1.795) 0.982 (3.090) 0.959 (2.158) 0.959 (2.158) 0.959 (2.158) 0.969 (3.095)
2.50 0.891 (2.181) 0.708 (3.090) 0.694 (2.545) 0.694 (2.545) 0.694 (2.545) 0.671 (2.032)
10 -2.50 0.977 (2.298) 0.988 (6.321) 1.000 (5.377) 1.000 (5.377) 1.000 (5.377) 0.999 (6.187)
-1.94 0.934 (2.052) 0.401 (5.869) 0.880 (5.377) 0.880 (5.377) 0.880 (5.377) 0.880 (5.085)
-1.39 0.940 (2.011) 1.000 (5.418) 1.000 (5.089) 1.000 (5.089) 1.000 (5.089) 1.000 (5.382)
-0.83 0.951 (2.011) 1.000 (4.966) 1.000 (5.171) 1.000 (5.171) 1.000 (5.171) 1.000 (5.186)
-0.28 0.931 (1.929) 0.997 (4.474) 1.000 (5.582) 1.000 (5.582) 1.000 (5.582) 0.979 (5.313)
0.28 0.944 (1.970) 1.000 (4.227) 1.000 (5.089) 1.000 (5.089) 1.000 (5.089) 0.957 (5.104)
0.83 0.945 (1.929) 0.943 (4.227) 1.000 (5.541) 1.000 (5.541) 1.000 (5.541) 0.973 (5.684)
1.39 0.947 (1.970) 1.000 (4.515) 1.000 (5.418) 1.000 (5.418) 1.000 (5.418) 0.983 (5.215)
1.94 0.955 (2.052) 0.997 (4.925) 1.000 (5.254) 1.000 (5.254) 1.000 (5.254) 0.995 (5.615)
2.50 0.982 (2.422) 0.997 (5.500) 1.000 (5.336) 1.000 (5.336) 1.000 (5.336) 0.998 (6.050)
B.3 Additional Experiments
B.3.1 Conditional Quantile Coverages
@@ Results
The data is visualized in Figure 9 in Appendix B. The resulting coverages are shown in Table 6. Our method provides (close to) nominal coverage for many combinations of $x$ and $\alpha$ but is less accurate for $x=0.2$ and extreme quantiles in general. This can be explained by the fact that the data at $x=0.2$ has almost no variation, while little data is available for extreme values, making these cases very challenging given only 100 data points.
Bootstrap Comparison
@@
B.3.2 Convergence Speed: Additional Results
Figures 6 and 7 depict the convergence speeds for the analysis in Section 4.2 for the two other learning rate schedules, type 1 and type 2, respectively. In this example, all results look qualitatively very similar.
@@ B.3.2 Convergence Speed: Additional Results
B.3.3 Benchmarks: Additional Results
Tables 8 and 9 summarize the results of our benchmark for $\alpha=0.5$ and $\alpha=0.9$, respectively for $n=25$. While both methods often have very close to nominal coverage or provide larger than nominal coverage, the bootstrap method fails in one case to provide a CI with close to nominal coverage. Despite having larger than nominal coverage, our MP approach consistently yields shorter or equally large CIs while often being twice as fast as the bootstrap.
@@ B.3.3 Benchmarks: Additional Results
B.3.4 Benchmarks: Comparison Using Fixed Time Budget
@@
Table 10: Coverage and interval lengths for MP and Bootstrap methods.
@@
B.4 Conditional Regression Posterior
Refer to caption Refer to caption Refer to caption
Results
B.5 Details on Data Generation
B.5.1 Data Generation for Diffusion Process
For the diffusion process data in Section B.4, we use the following data-generating process:
@@ B.5.1 Data Generation for Diffusion Process
B.5.2 Data Generation for Quantile Regression
@@
B.6 Computational Environment
All computations except for the pure TabPFN rollout were performed on a user PC with Intel(R) Core(TM) i7-8665U CPU @ 1.90GHz, 8 cores, 16 GB RAM, using Python 3.12.7. Run times of each experiment do not exceed 24 hours. For the TabPFN rollout, computations were done on a GeForce RTX 4070 Ti GPU, with every single experiment exceeding 1 hour of runtime. All experiments using TabPFN are run with tabpfn==2.0.5.
GPT 4o and GPT o3-mini-high are called via OpenAI’s UI. DeepSeek R1 is called from its UI at www.deepseek.com. All other methods are implemented from scratch by the authors.
Appendix C Further Discussion
Another open problem is joint posterior inference for a collection of parameters $\theta=\{\theta(x)\colon x\in\mathcal{X}\}$. Although our proposed inference procedure can be repeated for many values of $x$, the resulting posteriors are disconnected. For obtaining a full joint posterior $\Pi(\theta|\mathcal{D}_{n})$, the distribution of the features $x_{1:n}$ can no longer be ignored. Fong et al. (2023) proposed a general joint update of the PPDs for all values of $x$ simultaneously. However, this general update is intractable, and the heuristic simplifications proposed by Fong et al. (2023) are neither particularly simple nor theoretically justified. There are several potential ways forward. A simple practical solution would be to specify a joint distribution that combines the individual posteriors $P_{\infty,x_{1}},\dots,P_{\infty,x_{K}}$ in a plausible way; for example, using a multivariate Gaussian copula with covariance kernel depending on the distance between values $x_{i}\neq x_{j}$. We expect such a heuristic correction to work reasonably well in many applications. A more sophisticated alternative was recently proposed by Huk et al. (2024) and involves nonparametric estimation of the implicit dependence between PPDs by a nonparametric vine copula.
Appendix D LLM Proof-of-Concept
To substantiate our previous hypothesis, we provide a proof-of-concept by prompting different LLMs to construct a CDF based on $n=25$ using the setup of Falck et al. (2024). The prompt is given in the following:
@@ Appendix D LLM Proof-of-Concept
While Falck et al. (2024) investigated whether different LLMs possess the martingale property, we are interested in whether our approach can provide uncertainty for these models. To do so, we simulate i.i.d. data from a Gaussian distribution and provide it to the LLM (GPT-4o, GPT-o3-mini-high, DeepSeek R1) including a context that specifies how the data was generated. We then ask the LLM to generate CDF values $F(y_{i})$ for this distribution for a prespecified grid of $y_{i}$ values. Figure 10 shows the initial output CDF of the LLM as a red dashed line. The underlying true CDF is marked with a dashed-dotted green line, indicating that GPT-4o is unable to predict the CDF, while the others do so correctly. We then run $N=1000$ replications of forward loops of our Algorithm 1 to sample values $y_{(n+1):(n+N)}^{(b)}$ and plot the resulting CDF lines for $y_{1:(n+N)}^{(b)},b=1,\ldots,B$ in blue. These lines indicate the uncertainty of the CDF generated by the LLM. The average CDF resulting from these blue curves is plotted in yellow, indicating that all models have the martingale property (although GPT-4o is not correct in its prediction).
While the blue curves also nicely depict the epistemic uncertainty of the LLM for the predicted CDFs, it becomes clear that the provided method—as expected—cannot counteract biases in the LLM (as is the case for GPT 4o). For more advanced LLMs, however, our approach gives uncertainty quantification akin to what is investigated Requeima et al. (2024), but potentially much cheaper.
Posteriors for LLM-Generated Text
PFNs are in-context learners based on the transformer architecture, just as modern LLMs. LLMs are trained to approximate the PPD $p(y|x,\mathcal{D})$ of the next token $y$ given a context $x$ of previous tokens and the corpus $\mathcal{D}$ they were trained on. Given this similarity, the martingale posterior approach could potentially also be adapted to quantify uncertainty in LLM-generated text. There are, however, two major obstacles. First, simple PPD updates, as in (3), are impossible due to the complex relation between the next and the previous tokens. The complexity of such updates can likely only be matched by an equally strong LLM. Second, Proposition 3.1 and Proposition 3.2 fail due to the lack of exchangeability in the generated output tokens. Contraction results like Proposition 3.3, however, rely on only the martingale property and would remain valid, although the interpretation of the distributions $P_{M}$ for $M\to\infty$ is less clear. Investigating how and to what extent martingale posterior ideas transfer to language models is an exciting path for future research.